OpenStudy (anonymous):

x^2 + 2xy + (x^2 - 1/3x^3 - x^2y)dy/dx = 0 Solve this DE by finding a function of the form u = u(y), as an integrating factor.

OpenStudy (anonymous):

oh, okay, so what is a function u = u(y) such that both u and u' are present in the given equation

OpenStudy (anonymous):

x^2 and 1/3x^3 ?

OpenStudy (anonymous):

u = ? u' = ?

OpenStudy (anonymous):

u = 1/3 u ^3 u' = x^2 ?

OpenStudy (anonymous):

not u^3, and u' needs to be the first derivative of u

OpenStudy (anonymous):

sorry, i meant x!

OpenStudy (anonymous):

or would it be y?

OpenStudy (anonymous):

oh, okay no, though; try to replace as much of the given equation as possible by u and u'

OpenStudy (anonymous):

ok and then from there ..

OpenStudy (anonymous):

did you get something for u and u'?

OpenStudy (anonymous):

what do you mean did i get something for them? i thought you said just to sub them into the formula which gives me u' + 2xy + (u' - u + u'y)dy/dx = 0

OpenStudy (anonymous):

no, i said: oh, okay, so what is a function u = u(y) such that both u and u' are present in the given equation

OpenStudy (anonymous):

like...

OpenStudy (anonymous):

don't just substitute u and u' for x and y

OpenStudy (anonymous):

u is a function and u' is the derivative of that function

OpenStudy (anonymous):

i didn't? i subbed u = 1/3 x ^3 u' = x^2

OpenStudy (anonymous):

okay that's right - so sorry! - now, add more to u and u' so that you can reduce t he equation further by substitution

OpenStudy (anonymous):

you said try to replace as much as the given equatoin as possible with u and u' ..

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

like x^2*y and 2xy

OpenStudy (anonymous):

i'm supposed to add more to them? it's kind of unclear.

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

actually, start with -(x^2 + 2xy)dx = (x^2 - 1/3x^3 - x^2y)dy :-)

OpenStudy (anonymous):

ok.. so then dont start with the integrating factor subsitution?